I am stuck on correct answer for:
If x = 3 and y = -2, then 7x - 3y over 2x + y =
I believe answer is 27 over 4 but I need to verify my work. Can you show me the breakdown to get correct resolve?

Answer by anjulasahay(30)

I am stuck on correct answer for:
If x = 3 and y = -2, then 7x - 3y over 2x + y =
I believe answer is 27 over 4 but I need to verify my work. Can you show me the breakdown to get correct resolve?
ans: By putting the values of x and y given in the equation we get,
(7x-3y)/(2x+y)=(7*3-3*(-2))/(2*3+(-2))=(21+6)/(6-2)=27/4
your ans is correct.